OpenStudy (anonymous):

AB and BC form a right angle at point B. If A = (-3, -1) and B = (4, 4), what is the equation of BC?

OpenStudy (anonymous):

A. x + 3y = 16 B. 2x + y = 12 C. -7x − 5y = -48 D. 7x − 5y = 48

OpenStudy (anonymous):

what do u think